# Understanding Catastrophic Forgetting in Language Models via Implicit Inference

### v1

@@ 4.2 Effects of Reinforcement Learning from Human Feedback On Harmful Content Generation
The pretrained model might see examples of both capabilities, but RLHF will uniformly incentivize refusing to answer the question. Therefore, fine-tuning for safety can be framed as “catastrophically forgetting” how to answer the question.
Conjugate prompting to follow harmful instructions instead of refusing them.
In the context of our prior results, when we view language models as implicitly inferring the task, fine-tuning may be suppressing $\text{L}_{\text{ANSWER}}$ rather than forgetting it. Moreover, we expect that fine-tuning is primarily in English since preference data labelling is much more expensive and less diverse than unsupervised pretraining data (Hao, 2023). To test this claim, we test conjugate prompting to recover language model behavior before safety fine-tuning. Specifically, we test GPT-3.5 before and after fine-tuning for conversational dialogue  [^3] This corresponds to text-davinci-003 and gpt-3.5-turbo, respectively.. For our prompts, we sample $100$ instructions from AdvBench (Zou et al., 2023). We say that the model output reflects the ANSWER task if it attempts to answer the question, and otherwise reflects the REFUSE task if it is a refusal or an answer to a different question  [^4] We do not assess the correctness of the answer to the question since we are testing the success of the refusal mechanism. This is in line with research in this field such as Zou et al. (2023) and Wei et al. (2023a).. We provide concrete details for our setup and labelling guidelines in Appendix D.1 as well as examples of harmful instructions, model completions, and labels in Appendix D.2.
As can be seen in Table 2, fine-tuning takes the answer frequency from $92\%$ to $3\%$ in English. Similar to Section 4.1, we test whether we can apply conjugate prompting via language translation to recover the pretrained capability of answering harmful instructions. In line with our hypothesis, we find that the drop in performing the ANSWER task is always lower in non-English languages. For example, fine-tuning took the English ANSWER frequency from $92\%$ to $3\%$ while it took the Malayalam ANSWER frequency from $71\%$ to $65\%$. Therefore, we claim that $\text{L}_{\text{ANSWER}}$ is not forgotten, and conjugate prompting can recover this capability. We note that the brittleness of safety-training as well as conjugate prompting transformation functions have been concurrently documented by Wei et al. (2023a) in their comprehensive and impressive analysis of jailbreaking attacks.
Table 2: Measuring toxic generation vs refusal. We report the frequency that the model faithfully attempts to follow the harmful instruction vs the frequency it refuses or answers an unrelated question. We compare ChatGPT (gpt-3.5-turbo) against GPT-3.5 that hasn’t undergone safety fine-tuning for deployment (text-davinci-003). Each accuracy is taken over 100 harmful instructions from AdvBench ((Zou et al., 2023)). Every non-English language has a lower pretrained answer frequency and a lower frequency change than English.
Language GPT-3.5 Answer Freq ChatGPT Answer Freq Change in Frequency
@@ 5 Related Work
Understanding in-context learning. There has been a recent line of work on understanding how pretrained transformers perform in-context learning of simple functions. Garg et al. (2023); Li et al. (2023) study which classes can be in-context learnt,  Chan et al. (2022); Kirsch et al. (2022) study the conditions where in-context learning emerges, and Akyürek et al. (2022); von Oswald et al. (2022); Dai et al. (2023) focus on the exact in-context learning algorithm implemented in transformers. Inspired by these works, we focus on understanding in-context learning in the context of fine-tuning.
Another line of work focuses on how transformers implicitly determine which task to perform, with Xie et al. (2021) hypothesizing that next-token prediction task of pretraining can involve implicit bayesian inference; Min et al. (2022b); Wei et al. (2023b); Tamkin et al. (2022) construct experimental setups to probe how the prompts affect what task the model is inferring. Our work studies the same idea of task inference but builds on this work to first characterize the effect of fine-tuning and then intervene via conjugate prompting to switch between fine-tuned and pretrained behavior.
Fine-tuning pretrained language models. There is a large body of work on fine-tuning language models in a manner that preserves performance (Raffel et al., 2020; Arivazhagan et al., 2019; Gao et al., 2021), generalizes slightly out-of-distribution (Wei et al., 2022; Sanh et al., 2022; Min et al., 2022a), and aligns with human usage/values (Christiano et al., 2023; Stiennon et al., 2022; Bai et al., 2022; Ziegler et al., 2020; Ouyang et al., 2022; Mishra et al., 2022; Chung et al., 2022).
Catastrophic forgetting and continual learning. The general phenomenon of catastrophic forgetting where training on new tasks degrades performance on old tasks has been widely reported and studied (McCloskey & Cohen, 1989; Goodfellow et al., 2015; Kemker et al., 2017). There have been many attempts to address this problem, in what is customarily referred to as continual learning, via regularization or data replay (Kirkpatrick et al., 2017; Parisi et al., 2019; Peng & Risteski, 2022). In this work, we focus on the setting of fine-tuning in LLMs where we find more structure in the exact nature of catastrophic forgetting. This allows us to devise a prompting strategy to recover pretraining capabilities with no change to the actual fine-tuning procedure.
Prompting in different languages. Prior works have studied the multilingual capabilities of language models (Shi et al., 2022; Ahuja et al., 2023; Lin et al., 2022). The general finding is that transformers will best complete tasks in English with varying levels of performance drops in other languages. In this work, we focus on the disparity between pretraining and fine-tuning process across different languages (Hao, 2023).
Adversarial Attacks. Prior work and tweets have studied how to “jailbreak” LLMs to elicit undesirable content (Shin et al., 2020; Guo et al., 2021; Carlini et al., 2023; Zou et al., 2023). Specific instantiations of our conjugate prompting framework are closely related to some prior works. Ippolito et al. (2022) utilizes style transfer to circumvent hard-coded defenses for memorization-based attacks. In recent concurrent work, Wei et al. (2023a) report a similar finding of the possibility of jailbreaking via translation. We hope our work provides a unified perspective on such findings, along with a possible explanation via our hypothesis of shifting implicit inference.
Multi-task learning and meta-learning. The objective of learning to solve multiple tasks simultaneously falls under the broad umbrella of meta-learning (Finn et al., 2017; Kirsch & Schmidhuber, 2022; Andrychowicz et al., 2016) and multi-task learning (Evgeniou & Pontil, 2004; Radford et al., 2019). Within this literature, the work on  (Yin et al., 2020) provides an algorithm that allows for control over whether meta-learners utilize solutions to known tasks or generalize to new tasks. This is similar in objective to conjugate prompting but we focus only on manipulating the input rather than modifying the fine-tuning procedure.
@@ 6 Discussion and future work
We find that the apparent catastrophic effects of fine-tuning may be explained as shifting implicit task inference. We demonstrate that for fine-tuned models, transforming prompts to look dissimilar to the fine-tuning data via conjugate prompting can recover pretrained capabilities.
Contrary to heuristic prompt engineering, we provide a principled mechanism by which our strategy can recover pretraining performance. We acknowledge that using the pretrained model is a more direct way to obtain pretrained behavior when desired, but in the increasingly common blackbox API setting, we do not have access to all stages of model training (such as LLaMa-2 (Touvron et al., 2023b) and Claude). As such, conjugate prompting is also a warning that simply restricting access to only safety-finetuned models is not secure—we can still extract the pretrained behavior.
More than the immediate utility of conjugate prompting, we hope that our analysis of fine-tuning brings us closer to principled adaptation of pretrained models to applications of interest. We believe that our inference hypothesis opens up a number of interesting questions in terms of whether transformers explicitly execute task inference. A more mechanistic understanding of our phenomenon is still missing—are there sub-networks that perform task inference and capabilities, and could we manipulate these networks directly? Finally, it would also be interesting to develop better fine-tuning procedures, perhaps inspired by the continual learning and meta-learning literature. Better fine-tuning methods accompanied by a principled understanding would open up new avenues to guide task inference and leverage transformer capabilities, providing robust and steerable methods for deployment.
Broader Impact.
@@
7 Limitations
We acknowledge that the controlled nature of our hypothesis test comes at the cost of evaluating accuracy on more common NLP benchmarks. The specific instantiation of conjugate prompting through language translation may not be perfect since it relies on third-party services (e.g. Google Translate), is of lower quality in low-resource languages, and may not preserve the task for problems that rely on the language (e.g. famous quotes). Moreover, conjugate prompting requires domain-specific understanding of the tasks in the fine-tuning data and the task of interest.
@@ 8 Acknowledgements
We gratefully acknowledge the support of Apple and the AI2050 program at Schmidt Futures. JMS was supported by the NSF Graduate Research Fellowship. We thank Huan Zhang for providing compute for the linear regression experiments and Sang Michael Xie, Andrej Ristseki, Daniel Fried, and Jacob Steinhardt for helpful feedback in earlier stages of this work.
Appendix A Bayes Optimal Estimator for Mixture Distribution
A.1 Derivation
@@
Appendix B Regression Experiment Details
B.1 Change in Loss vs Likelihood under Fine-tuning Distribution
@@
B.2 Conjugate prompting for more $\alpha$ and $\gamma$
@@
B.3 Ridge regression is learnt before discrete regression on the discrete distribution
@@
B.4 Fine-Tuning on different mixtures
@@
B.5 Hyperparameters

### v2

@@ 4.3 Effects of RLHF On Harmful Content Generation
Conjugate prompting to follow harmful instructions.
Fine-tuning may be suppressing $\text{L}_{\text{ANSWER}}$ rather than forgetting it. Since preference data is more expensive and less diverse than pretraining data (Hao, 2023), we expect that fine-tuning is primarily in English, and we test conjugate prompting to recover behavior before safety fine-tuning. Specifically, we test GPT-3.5 before (gpt-3.5-turbo) and after (text-davinci-003) fine-tuning for conversational dialogue. For our prompts, we sample $100$ instructions from AdvBench (Zou et al., 2023). We say that the model output reflects the ANSWER task if it attempts to answer the question, and otherwise reflects the REFUSE task if it is a refusal or an answer to a different question (evaluation details in Appendix F.1, examples in Appendix F.2).
In line with our hypothesis, we find that the drop in the ANSWER task is always lower in non-English languages. For example, fine-tuning took the English ANSWER frequency from $92\%$ to $3\%$ while it took the Malayalam frequency from $71\%$ to $65\%$. Therefore, we claim that conjugate prompting can partially recover the capability of harmful instruction following. We note that the brittleness of safety-training as well as transformation functions have been concurrently documented by Wei et al. (2023a) in their comprehensive and impressive analysis of jailbreaking attacks.
Table 2: Measuring accuracy on XNLI after code fine-tuning. We report XNLI accuracy over 2490 test samples. There is a drop in English accuracy after code fine-tuning. For Spanish, French, and German, the accuracy barely changes or slightly increases, performing best in Spanish.
Language LLaMa-2 Acc Code LLaMa Acc Drop
English
French
Spanish
German
44.26 %
33.53 %
38.11 %
34.50 %
35.90 %
34.98 %
38.88 %
33.49 %
8.36 %
-1.45 %
-0.77 %
1.01 %
Table 3: Measuring toxic generation vs refusal. We measure whether the model attempts to follow a harmful instruction. We compare ChatGPT against GPT-3.5 without safety fine-tuning. Each cell is taken over 100 harmful instructions. Every non-English language has a lower pretrained answer frequency and a lower frequency change than English.
Language GPT-3.5 Answer ChatGPT Answer Drop
@@ 5 Related Work
Additional related work on ICL, fine-tuning, and multilingual NLP can be found in Appendix A.
Catastrophic forgetting and continual learning. Catastrophic forgetting has been widely studied (McCloskey & Cohen, 1989; Goodfellow et al., 2015; Kemker et al., 2017) with several works assessing its prevalence in modern settings (Ramasesh et al., 2022; Luo et al., 2023; Wang et al., 2023). There have been many attempts to address this through continual learning algorithms and data replay (Kirkpatrick et al., 2017; Parisi et al., 2019; Peng & Risteski, 2022). We focus on leveraging extra problem structure in the LLM setting to devise our prompting strategy.
Multi-task learning and meta-learning. Learning to solve multiple tasks falls under meta-learning (Finn et al., 2017; Kirsch & Schmidhuber, 2022; Andrychowicz et al., 2016) and multi-task learning (Evgeniou & Pontil, 2004; Radford et al., 2019). For example, (Yin et al., 2020) provides a training algorithm to control whether meta-learners perform known tasks or generalize to new tasks. Unlike prior work, we focus on manipulating the input rather than modifying training.
Adversarial Attacks. Prior work/tweets have studied how to “jailbreak” LLMs to elicit undesirable content (Shin et al., 2020; Guo et al., 2021; Carlini et al., 2023; Zou et al., 2023). Instances of our framework have been studied, such as attacks via translation (Wei et al., 2023a) and style transfer to elicit memorized content (Ippolito et al., 2022). We hope to provide a unified perspective.
@@ 6 Discussion and future work
We find that the catastrophic effects of fine-tuning may be explained as shifting task inference and that transforming prompts further from the fine-tuning data can recover pretrained capabilities. This becomes important in the increasingly common blackbox API setting (i.e. ChatGPT, Claude), where conjugate prompting also warns that restricting access to safety fine-tuned models is not secure.
More than immediate utility, we hope our analysis brings us closer to principled adaptation of pretrained models. Our inference hypothesis opens up interesting questions in terms of whether transformers explicitly execute task inference through sub-networks that we could manipulate directly. Finally, better fine-tuning methods accompanied by a principled understanding could open up robust methods to guide task inference and leverage transformer capabilities for deployment.
Limitations.
Translation is not perfect due to third-party services, low-resource languages, and contextual knowledge. Conjugate prompting requires knowledge of training data and deployment tasks. There is scope for larger evaluation testing relationships involving data, model size, and tasks.
7 Ethics Statement
@@ 8 Acknowledgements
We thank Huan Zhang for providing compute for the linear regression experiments and Sang Michael Xie, Andrej Ristseki, Daniel Fried, Jacob Steinhardt, and Ankit Bisain for helpful feedback in earlier stages of this work.
We gratefully acknowledge the support of Apple and the AI2050 program at Schmidt Futures. JMS was supported by the NSF Graduate Research Fellowship. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE2140739. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors(s) and do not necessarily reflect the views of the National Science Foundation.
Appendix A Additional Related Work
Understanding in-context learning. There has been a recent line of work on understanding how pretrained transformers perform in-context learning of simple functions. Garg et al. (2023); Li et al. (2023) study which classes can be in-context learnt,  Chan et al. (2022); Kirsch et al. (2022) study the conditions where in-context learning emerges, and Akyürek et al. (2022); von Oswald et al. (2022); Dai et al. (2023) focus on the exact in-context learning algorithm implemented in transformers. Inspired by these works, we focus on understanding in-context learning in the context of fine-tuning. Another line of work focuses on how transformers implicitly determine which task to perform, with Xie et al. (2021) hypothesizing that next-token prediction task of pretraining can involve implicit bayesian inference; Min et al. (2022b); Wei et al. (2023b); Tamkin et al. (2022) construct experimental setups to probe how the prompts affect what task the model is inferring. Our work studies the same idea of task inference but builds on this work to first characterize the effect of fine-tuning and then intervene via conjugate prompting to switch between fine-tuned and pretrained behavior.
Fine-tuning pretrained language models. There is a large body of work on fine-tuning language models in a manner that preserves performance (Raffel et al., 2020; Arivazhagan et al., 2019; Gao et al., 2021b), generalizes slightly out-of-distribution (Wei et al., 2022; Sanh et al., 2022; Min et al., 2022a), and aligns with human usage/values (Christiano et al., 2023; Stiennon et al., 2022; Bai et al., 2022; Ziegler et al., 2020; Ouyang et al., 2022; Mishra et al., 2022; Chung et al., 2022). Other works have tried to build a mechanistic understanding for how fine-tuning alters (or does not alter) pretrained models (Lubana et al., 2023; Jain et al., 2023).
Prompting in different languages. Prior works have found that models will best complete tasks in English with performance drops in other languages (Shi et al., 2022; Ahuja et al., 2023; Lin et al., 2022). We highlight the disparity of this phenomenon between pretraining and fine-tuning.
Appendix B Bayes Optimal Estimator for Mixture Distribution
B.1 Derivation
@@
Appendix C Regression Experiment Details
C.1 Change in Loss vs Likelihood under Fine-tuning Distribution
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
C.2 Conjugate prompting for more $\alpha$ and $\gamma$
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
C.3 Ridge regression is learnt before discrete regression on the discrete distribution
@@
C.4 Fine-Tuning on different mixtures
@@
Since this achieves better tradeoff between the two losses, we know that standard fine-tuning is necessarily catastrophic: if it wasn’t, it would achieve the tradeoff we see in this section.
Refer to caption Refer to caption
@@
C.5 Task Ambiguity
In this section, we quantify the ambiguity present in the original pretraining task. Specifically, we consider whether mixture regression can accurately distinguish between discrete tasks from $\mathcal{D}_{\text{disc}}$ and continuous tasks from $\mathcal{D}_{\text{cont}}$. We visualize the Bayes-optimal $g(X,y)$ for $\alpha\in\{0.2,0.5,0.8\}$ and exemplar counts $\{5,10,15\}$ in Figure 13.
To demonstrate the pretrained model can distinguish the continuous distribution from the discrete distribution, we plot the continuous loss of a model pretrained on $\alpha=0.5$ in Figure 14. We find that the model performs much closer to ridge regression than a fine-tuned model (Figure 10, middle).
We find that the true posterior and the pretrained model can relatively easily discern between continuous and discrete tasks, especially for exemplar count $10$ and higher; this demonstrates how there is little ambiguity in the original task. Regardless, the fine-tuned model does not perform ridge regression for prompts from $\mathcal{D}_{\text{cont}}$ after fine-tuning. We aim to investigate whether the model has forgotten how to do ridge regression, or whether it has changed its internal posterior to perform discrete regression for these tasks. Conjugate prompting supports our hypothesis that fine-tuning is changing task inference rather than only degrading pre-existing capabilities.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: True posterior $g(X,y)$ for pretraining distribution. We plot the distribution of the true posterior of the discrete distribution $g(X,y)$ for $\alpha=\{0.2,0.5,0.8\}$ when sampling from the continuous distribution $\mathcal{D}_{\text{cont}}$ and discrete distribution $\mathcal{D}_{\text{disc}}$ for $5,10,15$ exemplars. We find that an optimal pretrained model can effectively infer whether the task is from $\mathcal{D}_{\text{cont}}$ or $\mathcal{D}_{\text{disc}}$, especially for $\geq 10$ samples. We note that the posterior for $\mathcal{D}_{\text{cont}}$ is the complement (horizontal reflection) of these plots. The violin plots are constructed by taking 2048 samples from the respective distribution and cutting any density estimate outside the support.
Figure 14: Model pretrained on $\alpha=0.5$ mixture. We find that the pretrained model for $\alpha=0.5$ performs close to ridge regression, showing how the pretrained model can effectively distinguish between the continuous and discrete distributions much better than the model fine-tuned on $\mathcal{D}_{\text{disc}}$ after being pretrained on $\alpha=0.5$ (Figure 6).
C.6 Reproduction for Larger Models
We are interested in seeing whether our experiments are consistent across model size. In our main experiments, we use a GPT-2 model with embedding dimension 256, 12 layers, 8 heads, and 22.4 million parameters. In this section, we experiment with a larger model of embedding dimension 384, 18 layers, 12 heads, and 51.3 million parameters, which is double the original parameter count.
In Figure 15, we plot the loss when fine-tuning our larger model for 400 steps on the discrete distribution after pretraining for 5000 steps on $\alpha=0.2$. We find that catastrophic forgetting still exists as there is a sudden drop in loss on the discrete distribution and a sudden spike in loss on the continuous distribution. We also see that the drop is slightly smaller as the base model is larger, which is expected since the base discrete loss is smaller with a stronger model.
Figure 15: Fine-tuning hurts continuous loss for larger model. We train an $\alpha=0.2$ large transformer with $64$ discrete tasks for $5000$ steps and fine-tune for $400$ steps on $\mathcal{D}_{\text{disc}}$ (highlighted). The discrete loss rapidly decreases, while the continuous loss rapidly increases.
We now test conjugate prompting for the larger model after pretraining on $\alpha=\{0.5\}$ for 5000 steps (Figure 16). We find that conjugate prompting consistently helps. Similar to our other experiments, it helps the most at low exemplar counts and more when the prior is already more biased to fine-tuning tasks. The benefits of conjugate prompting seem similar across scale, though the fine-tuned models for standard continuous prompts seems slightly worse for the larger model, potentially due to stronger base performance on the discrete distribution. We also quantify these results in Table 4 for $\alpha\in\{0.2,0.5,0.8\}$. At a high level, we find that both larger and smaller models forget at similar rates, and conjugate prompting helps both models.
Refer to caption Refer to caption Refer to caption
Figure 16: Conjugate prompting for larger models. We take the larger transformers pretrained over $\mathcal{D}_{\text{mix}}$ for $\alpha\in\{0.2,0.5,0.8\}$ for $5000$ steps and fine-tuned over $\mathcal{D}_{\text{disc}}$ for $400$ steps. We evaluate their loss on the continuous distribution where they under-perform ridge regression. Conjugate prompting with scale factor $\gamma\in\{1.5,2.0\}$ recovers the pretrained solution of ridge regression, especially on lower sample counts with more ambiguity.
Table 4: Measuring forgetting over model scale. We quantify the level of forgetting and the success of conjugate prompting across model scale. To do this, we take the loss of the model at 3 stages: before fine-tuning, after fine-tuning, and after conjugate prompting. We find that the drop is larger for the 22.4M model for $\alpha=0.5,0.8$ and the drop is larger for the the 51.3M model for $\alpha=0.2$. The losses are averaged over $4096$ sequence samples with 00 to $20$ exemplars, similar to conjugate prompting plots such as Figure 6.
Mixture $\alpha$ Parameter Count Before FT After FT FT Drop Conjugate Prompting $\gamma=1.5$ Drop ($\downarrow$) Conjugate Prompting $\gamma=2.0$ Drop ($\downarrow$)
$\alpha=0.2$
22.4M
51.3M
$0.625$
$0.625$
$0.867$
$0.892$
$0.242$
$0.267$
$0.163$
$0.181$
$0.170$
$0.177$
$\alpha=0.5$
22.4M
51.3M
$0.641$
$0.648$
$0.882$
$0.876$
$0.241$
$0.228$
$0.161$
$0.108$
$0.170$
$0.094$
$\alpha=0.8$
22.4M
51.3M
$0.662$
$0.653$
$0.861$
$0.833$
$0.199$
$0.180$
$0.091$
$0.097$
$0.086$
$0.101$
C.7 Reproduction for Different Dataset Size
We are interested in seeing whether our experiments are consistent across dataset size. In our main experiments, we pretrain the model on 5000 steps. In this section, we experiment with models pretrained on their corresponding mixtures for less steps (2000) and more steps (10000).
In Figures 17 and 18, we plot the loss when fine-tuning a model for 400 steps on the discrete distribution after pretraining for 2000 steps or 10000 steps on $\alpha=0.2$. We find that forgetting still exists as there is a sudden drop in loss on the discrete distribution and a sudden spike on the continuous distribution. We also see that the drops and spikes are slightly smaller as the model is pretrained for longer, which is expected since the base discrete loss is smaller with more data.
Figure 17: Fine-tuning hurts continuous loss for 2000 pretraining steps. We train an $\alpha=0.2$ transformer with $64$ discrete tasks for $2000$ steps and fine-tune for $400$ steps on $\mathcal{D}_{\text{disc}}$ (highlighted). The discrete loss rapidly decreases, while the continuous loss rapidly increases.
Figure 18: Fine-tuning hurts continuous loss for 10000 pretraining steps. We train an $\alpha=0.2$ transformer with $64$ discrete tasks for $10000$ steps and fine-tune for $400$ steps on $\mathcal{D}_{\text{disc}}$ (highlighted). The discrete loss rapidly decreases, while the continuous loss rapidly increases.
We now test conjugate prompting for fine-tuned models after pretraining on $\alpha=\{0.2,0.5,0.8\}$ for 2000 steps (Figure 19) or 10000 steps (Figure 20). We find that conjugate prompting consistently helps. Similar to our other experiments, it helps the most at low exemplar counts and more when the prior is already more biased to fine-tuning tasks. The benefits of conjugate prompting seem similar across scale, presumably since fine-tuning takes base models to similar functions.
Refer to caption Refer to caption Refer to caption
Figure 19: Conjugate prompting for 2000 pretraining steps. We take transformers pretrained over $\mathcal{D}_{\text{mix}}$ for $\alpha\in\{0.2,0.5,0.8\}$ for $2000$ steps and fine-tuned over $\mathcal{D}_{\text{disc}}$ for $400$ steps. We evaluate their loss on the continuous distribution where they under-perform ridge regression. Conjugate prompting with scale factor $\gamma\in\{1.5,2.0\}$ recovers the pretrained solution of ridge regression, especially on lower sample counts with more ambiguity.
Refer to caption Refer to caption Refer to caption
Figure 20: Conjugate prompting for 10000 pretraining steps. We take transformers pretrained over $\mathcal{D}_{\text{mix}}$ for $\alpha\in\{0.2,0.5,0.8\}$ for $10000$ steps and fine-tuned over $\mathcal{D}_{\text{disc}}$ for $400$ steps. We evaluate their loss on the continuous distribution where they under-perform ridge regression. Conjugate prompting with scale factor $\gamma\in\{1.5,2.0\}$ recovers the pretrained solution of ridge regression, especially on lower sample counts with more ambiguity.
C.8 Hyperparameters